OpenStudy (anonymous):

13a-15b-a+2b= step by step please thanks

OpenStudy (angela210793):

=12a-13b 13-1=12 -15+2=-13 just like with no.s

OpenStudy (anonymous):

oh ok, so it was just combining like terms basically?

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

oh lol im so dumb thanks so much (:

OpenStudy (angela210793):

lol...uw :)

OpenStudy (anonymous):

13a-15b-a+2b =13a-a-15b+2b =12a-13b (answer)